A unitörm distribution has a mean oft and Variance 4/3. Find P(Xe0).

1
Expert's answer
2021-12-07T09:56:51-0500
"\mu=\dfrac{A+B}{2}=1"

"\sigma^2=\dfrac{(B-A)^2}{12}=\dfrac{4}{3}"

"A+B=2"

"B-A=4"

"A=-1, B=3"

"f(x)= \begin{cases}
 \dfrac{1}{3-(-1)} &-1\leq x\leq 3 \\
 0 & otherwise
\end{cases}"

"P(X<0)=\displaystyle\int_{-1}^{0}\dfrac{1}{4}dx=[\dfrac{1}{4}x]\begin{matrix}
 0 \\
 -1
\end{matrix}=\dfrac{1}{4}"

"P(X<0)=0.25"